# Dijets with a large rapidity separation at the next-to-leading BFKL for search of large extra dimension gravity at colliders

### v3

@@ Abstract
Search for the gravity with large extra dimensions at collider energies is considered in the trans-Planckian eikonal regime, i. e., when $\mbox{$\sqrt{\hat{s}}$}\gg\mbox{$M_{D}$}\gg\sqrt{-\mbox{$\hat{t}$}}$. Here $\hat{s}$ and $\hat{t}$ are the Mandelstam variables of colliding parton-parton system and $M_{D}$ is the Planck mass scale in the space-time with compactified $n_{D}$ extra dimensions. A relevant observable for this regime may be the cross section of high-mass ($\mbox{$M_{jj}$}\sim\mbox{$\sqrt{\hat{s}}$}\gg\mbox{$M_{D}$}$) dijet production with a large rapidity separation. Then the standard model (SM) background should be calculated within the next-to-leading logarithmic (NLL) approximation of Lipatov-Fadin-Kuraev-Balitsky (BFKL) formalism of quantum chromodynamics (QCD) suitable for $\sqrt{\hat{s}}\gg\sqrt{-\hat{t}}\gg\Lambda_{\mathrm{QCD}}$. In this work the signal of the large extra dimension gravity as well as the NLL BFKL QCD background are estimated for the high-luminosity Large Hadron Collider (HL-LHC) and future colliders such as FCCpp and CEPC-SppC.
@@ I INTRODUCTION
Recently, it was shown that cross section of dijet production with a large rapidity separation at the LHC [8] can be described by the NLL BFKL evolution [10], while DGLAP prediction significantly overestimates the data. Therefore, searches for new physics beyond the SM at the high-luminosity Large Hadron Collider (HL-LHC) [22] and future colliders, such as FCCpp [23] and CEPC-SppC [24], by dijets with a large rapidity separation should take into account new BFKL evolution dynamics for background estimates.
@@ II Gravity in space-time with large compactified extra dimensions
The concept behind the ADD model which explains the observed weakness of gravity relative to other fundamental forces is simple. The graviton is expected to propagate through the usual $4$-dimensional space-time as well as $n_{D}$ extra spatial dimensions, whereas the other fundamental fields are confined to the $4$-dimensional space-time submanifold. The gravitational potential between two masses in the presence of $n_{D}$ extra dimensions behaves as:
@@
where $M_{D}$ is the fundamental parameter that sets the characteristic scale of gravity in the $4+n_{D}$-dimensional space-time. If the $n_{D}$ extra dimensions are compactified with size $R$, this modified gravitational potential behavior can be observed only at short distances where $r\ll R$. At large distances $r\gg R$, the gravitational potential is expected to follow the ordinary $1/r$ behavior:
$V(r)\sim\frac{1}{M_{D}^{2+n_{D}}R^{n_{D}}}\frac{1}{r}$ (2)
This immediately shows that the large magnitude of the Planck mass $M_{\mathrm{Pl}}$ is determined by the volume of the extra dimensions, such that $M_{\mathrm{Pl}}=M_{D}^{1+n_{D}/2}R^{n_{D}/2}$. Assuming that the gravity scale $M_{D}$ is on the order of the electroweak scale, $m_{\mathrm{EW}}$, the required size $R$ of the extra dimensions can be determined as a function of their number $n_{D}$:
$R\sim 10^{\frac{30}{n_{D}}-17}\mathrm{cm}\times\bigg(\frac{1\mathrm{TeV}}{m_{\mathrm{EW}}}\bigg)^{1+\frac{2}{n_{D}}}$ (3)
@@
The work by GRW Ref. [27] explores the possibility of parton scattering through multiple KK graviton exchanges in the trans-Planckian regime using the eikonal approximation, where $\mbox{$\sqrt{\hat{s}}$}\gg\mbox{$M_{D}$}\gg\sqrt{-\mbox{$\hat{t}$}}$. In this kinematic regime, the gravity signal is expected to manifest itself as an excess in the production of dijets with jets which are widely separated in rapidity. This approach resums an infinite series of graviton exchanges, valid when the momentum transfer is small compared to the center-of-mass energy. The summation of an infinite series of successive gluon exchanges results in the following expression for the parton-parton cross section:
$\frac{d\hat{\sigma}_{\mathrm{eik}}}{d\Delta y}=\frac{\pi b_{c}^{4}\hat{s}e^{\Delta y}}{(1+e^{\Delta y})^{2}}\big|F_{D}(x)|^{2}$ (4)
where $b_{c}$ end $F_{D}(x)$ are defined by formulas (18) and (20) of Ref. [27]; $x=b_{c}\sqrt{\hat{s}}/\sqrt{1+e^{\Delta y}}$; and $\Delta y$ is the separation in rapidity between the two jets of a dijet. In the trans-Planckian ($\sqrt{\hat{s}}/M_{D}\gg 1$) eikonal ($\hat{t}/\hat{s}\ll 1$) regime $\sqrt{\hat{s}}$ is equal to jet-jet invariant mass $M_{jj}=\sqrt{\hat{s}}$, and $\hat{t}/\hat{s}=1/(1+e^{\Delta y})$. Using the parton-parton cross section (4), one can convolute it with parton distribution functions (PDFs) to obtain the dijet production cross section in $pp$ collisions due to multiple graviton exchanges. It is important to note that the cross section (4) is valid only in the regime where $\hat{t}/\hat{s}\ll 1$, corresponding to large rapidity intervals ($\Delta y\gg 1$). How this expression matches the region of small $\Delta y$ remains basically unclear.
@@ III QCD background estimation
Recent results from the CMS experiment at the LHC report the first measurement of the cross section for Mueller-Navelet (MN) dijet production at TeV-scale energies [8]. MN dijets refer to the pair of jets in an event that have the largest separation in rapidity among all jets with transverse momentum above a given threshold $p_{\perp\min}$ [37]. The CMS collaboration also provided comparisons of their measurements with theoretical calculations that incorporate leading order (LO) and next-to-leading order (NLO) QCD matrix elements, enhanced by leading logarithmic (LL) DGLAP evolution of the parton shower. The comparison revealed that theoretical models using DGLAP evolution significantly overestimate the measured dijet production cross sections, by up to a factor of six in the region of large rapidity separation $\Delta y$. Noticeable deviations of DGLAP-based calculations from the measurements begin to appear for rapidity separations $\mbox{$\Delta y$}\gtrsim 4$, indicating that the theoretical models based on the DGLAP evolution, become less accurate in this large rapidity interval region.
The accuracy of LL BFKL-based calculations is generally insufficient, as the LL approximation tends to predict stronger effects than those observed in experimental measurements. This overestimation arises due to limitations inherent in the LL BFKL approach, such as neglecting certain higher-order corrections like running of the strong coupling and kinematical constraints such as energy conservation. Consequently, next-to-leading logarithmic (NLL) corrections are necessary to achieve a more accurate and realistic description of the data [20, 21]. Calculations performed in Ref. [10], based on the NLL BFKL evolution, demonstrate good agreement with the measurement data in the region $\mbox{$\Delta y$}>4$. This provides new evidence supporting the possible manifestation of BFKL evolution effects at TeV energy scales and suggests that the QCD background is more accurately described by the NLL BFKL evolution in the regime of large $\Delta y$.
The NLL BFKL calculations rely on three key components: the process-independent BFKL Green’s function, process-dependent impact factors, and a procedure to resolve the QCD renormalization scale ambiguity. The detailed calculations for the BFKL Green’s function and impact factors for MN dijets can be found in Refs [38, 39]. The approach that systematically eliminates scale-setting ambiguities was developed by Brodsky, Fadin, Kim, Lipatov and Pivovarov (BFKLP) [40]. This method essentially generalizes the Brodsky, Lepage and Mackenzie (BLM) optimal scale-setting procedure [41] to the non-Abelian case.
@@ IV Definition of observable
This study examines $pp$ collisions at energies of $\mbox{$\sqrt{s}$}=13$, $40$ and $100$ TeV, spanning from current LHC energies up to those expected for FCCpp and CEPC-SppC. Although the ADD signal to QCD background ratio is expected to improve with increasing $M_{jj\mathrm{min}}$, the cross sections diminish significantly at large $M_{jj\mathrm{min}}$. Thus, the high dijet mass cuts are selected based on the projected integrated luminosities of the HL-LHC, FCCpp and CEPC-SppC. Specifically, for $\mbox{$\sqrt{s}$}=13$ TeV we consider $\mbox{$M_{jj\mathrm{min}}$}=6$, and $9$ TeV; for $\mbox{$\sqrt{s}$}=40$ TeV, $\mbox{$M_{jj\mathrm{min}}$}=9$, and $30$ TeV; and for $\mbox{$\sqrt{s}$}=100$ TeV, $\mbox{$M_{jj\mathrm{min}}$}=30$, and $70$ TeV.
@@
$M_{jj}^{2}=2p_{\perp 1}p_{\perp 2}(\cosh{(\mbox{$\Delta y$})}-\cos(\Delta\phi))$ (5)
where $\Delta\phi$ represents the difference in azimuthal angle between the two jets forming the dijet. From here we see that imposition high $M_{jj\mathrm{min}}$ selection on a dijet effectively select jets with $p_{\perp}$ higher than $p_{\perp\min}$. In the BFKL kinematic regime, $p_{\perp 1}$ can differ from $p_{\perp 2}$ due to transverse momentum diffusion. The effective minimum jet transverse momentum, $p_{\perp\min}^{\mathrm{eff}}$, is determined by $M_{jj\mathrm{min}}$, $\Delta y$ and $p_{\perp\max}=\mbox{$\sqrt{s}$}/(2\cosh(\max[-y_{\max}+\mbox{$\Delta y$},0]))$. For the selections analyzed in this study, the effective minimum jet $p_{\perp}$ in BFKL kinematics is shown in Fig 1a. It is observed that $\mbox{$p_{\perp\min}^{\mathrm{eff}}$}<\mbox{$p_{\perp\min}$}=20$ GeV only for the $\mbox{$\sqrt{s}$}=40$ TeV with $\mbox{$M_{jj}$}>9$ TeV selection, where $\mbox{$p_{\perp\min}^{\mathrm{eff}}$}\approx 18$ GeV for $\mbox{$\Delta y$}>6$. However, cases with a large imbalance, i.e. when the $p_{\perp}$ of one jet is much smaller than that of the other, lead to an oscillating integrand in the BFKL calculation, causing these configurations to be suppressed relative to those with $p_{\perp 1}\approx p_{\perp 2}$. When $p_{\perp 1}=p_{\perp 2}$, as in the DGLAP kinematics, the effective minimum transverse momentum, $p_{\perp\min}^{\mathrm{eff}}$, is determined by $M_{jj\mathrm{min}}$ and $\Delta y$. The corresponding $p_{\perp\min}^{\mathrm{eff}}$ values are presented in Fig 1b. It is evident that $p_{\perp\min}^{\mathrm{eff}}$ is significantly higher than $\mbox{$p_{\perp\min}$}=20$ GeV in this case.
@@
V Numerical calculations
@@
The QCD background within the DGLAP approximation is evaluated by either convolving LO QCD partonic cross sections with NNPDF31_lo_as0130 PDFs evolved through LL DGLAP evolution or using the pythia8 MC generator with the CP5 tune [47]. Originally pythia8 performed at LO+LL DGLAP accuracy as well. However, the CP5 tune updates it by using a NLO strong coupling constant and NNLO PDFs, along with rapidity ordering in the initial-state radiation. This rapidity ordering makes behavior of pythia8 predictions closer to the BFKL kinematics, improving the modeling of gluon emissions and parton evolution for high-energy collisions. The validity of these adjustments require careful theoretical consistency checks. We developed a custom pythia8 plugin to improve computational efficiency in MC simulations by biasing event generation toward high dijet masses at the parton level and correcting this bias with compensating weights to maintain accurate overall distributions. The plugin uses dijet mass at parton level, and its performance was validated against unweighted simulations. The applied exponential weight-bias function enhances computational efficiency of high dijet mass event generation while keeping weights bounded and uncertainties manageable. The form of the bias function is:
$\exp\left(\frac{\sqrt{\hat{s}}}{a}\right)-1+\frac{b}{a}$ (6)

### v4

# Dijets with a large rapidity separation in the next-to-leading order BFKL formalism for searches of large extra dimensions at colliders

@@ Abstract
Search for the gravity with large extra dimensions at collider energies is considered in the trans-Planckian eikonal regime, i.e., when $\mbox{$\sqrt{\hat{s}}$}\gg\mbox{$M_{D}$}\gg\sqrt{-\mbox{$\hat{t}$}}$. Here $\hat{s}$ and $\hat{t}$ are the Mandelstam variables of colliding parton-parton system and $M_{D}$ is the Planck mass scale in the space-time with compactified $n_{D}$ extra dimensions. A relevant observable for this regime may be the cross section of high-mass ($\mbox{$M_{jj}$}\sim\mbox{$\sqrt{\hat{s}}$}\gg\mbox{$M_{D}$}$) dijet production with a large rapidity separation. Then the standard model background should be calculated within the next-to-leading logarithmic (NLL) approximation of Lipatov-Fadin-Kuraev-Balitsky (BFKL) formalism of quantum chromodynamics (QCD) suitable for $\sqrt{\hat{s}}\gg\sqrt{-\hat{t}}\gg\Lambda_{\mathrm{QCD}}$. In this work the signal of the large extra dimension gravity as well as the NLL BFKL QCD background are estimated for the high-luminosity Large Hadron Collider (HL-LHC) and future colliders such as FCCpp and CEPC-SppC.
@@ I INTRODUCTION
Recently, it was shown that cross section of dijet production with a large rapidity separation at the LHC [8] can be described by the NLL BFKL evolution [10], while DGLAP prediction significantly overestimates the data. Therefore, searches for new physics beyond the SM at the high-luminosity Large Hadron Collider (HL-LHC) [22] and future colliders, such as FCCpp [23] and CEPC-SppC [24], by dijets with a large rapidity separation should take into account new BFKL evolution dynamics for SM background estimates.
@@ II Gravity in space-time with large compactified extra dimensions
The concept behind the ADD model which explains the observed weakness of gravity relative to other fundamental forces is the following [25]: the graviton is expected to propagate through the usual four-dimensional space-time as well as $n_{D}$ extra spatial dimensions, whereas the other fundamental fields are confined to the four-dimensional space-time submanifold. The gravitational potential between two masses in the presence of $n_{D}$ extra dimensions behaves as
@@
where $M_{D}$ is the fundamental parameter that sets the characteristic scale of gravity in the $4+n_{D}$-dimensional space-time. If the $n_{D}$ extra dimensions are compactified with size $R$, this modified gravitational potential behavior can be observed only at short distances where $r\ll R$. At large distances $r\gg R$, the gravitational potential is expected to follow the ordinary $1/r$ behavior
$V(r)\sim\frac{1}{M_{D}^{2+n_{D}}R^{n_{D}}}\frac{1}{r}.$ (2)
This immediately shows that the large magnitude of the Planck mass $M_{\mathrm{Pl}}$ is determined by the volume of the extra dimensions, such that $M_{\mathrm{Pl}}=M_{D}^{1+n_{D}/2}R^{n_{D}/2}$. Assuming that the gravity scale $M_{D}$ is on the order of the electroweak scale, $m_{\mathrm{EW}}$, the required size $R$ of the extra dimensions can be determined as a function of their number $n_{D}$
$R\sim 10^{\frac{30}{n_{D}}-17}\mathrm{cm}\times\bigg(\frac{1\mathrm{TeV}}{m_{\mathrm{EW}}}\bigg)^{1+\frac{2}{n_{D}}}.$ (3)
@@
The work by GRW Ref. [27] explores the possibility of parton scattering through multiple KK graviton exchanges in the trans-Planckian regime using the eikonal approximation, where $\mbox{$\sqrt{\hat{s}}$}\gg\mbox{$M_{D}$}\gg\sqrt{-\mbox{$\hat{t}$}}$. In this kinematic regime, the gravity signal is expected to manifest itself as an excess in the production of dijets with jets which are widely separated in rapidity. This approach resums an infinite series of graviton exchanges, valid when the momentum transfer is small compared to the center-of-mass energy. The summation of an infinite series of successive graviton exchanges results in the following expression for the parton-parton cross section:
$\frac{d\hat{\sigma}_{\mathrm{eik}}}{d\Delta y}=\frac{\pi b_{c}^{4}\hat{s}e^{\Delta y}}{(1+e^{\Delta y})^{2}}\big|F_{D}(x)|^{2},$ (4)
where $b_{c}$ is a length scale defining the upper boundary of classical gravity consideration. The function $F_{D}(x)$ is the essential component of the eikonal amplitude. The definitions of $b_{c}$ and $F_{D}(x)$ are given by Eqs. (18) and (20) of Ref. [27]; $x=b_{c}\sqrt{\hat{s}}/\sqrt{1+e^{\Delta y}}$; and $\Delta y$ is the separation in rapidity between the two jets of a dijet. In the trans-Planckian ($\sqrt{\hat{s}}/M_{D}\gg 1$) eikonal ($\hat{t}/\hat{s}\ll 1$) regime $\sqrt{\hat{s}}$ is equal to jet-jet invariant mass $M_{jj}=\sqrt{\hat{s}}$, and $\hat{t}/\hat{s}=1/(1+e^{\Delta y})$. Using the parton-parton cross section equation (4), one can convolute it with parton distribution functions (PDFs) to obtain the dijet production cross section in $pp$ collisions due to multiple graviton exchanges. It is important to note that the cross section equation (4) is valid only in the regime where $\hat{t}/\hat{s}\ll 1$, corresponding to large rapidity intervals ($\Delta y\gg 1$). How this expression matches the region of small $\Delta y$ remains basically unclear.
@@ III QCD background estimation
Recent results from the CMS experiment at the LHC report the first measurement of the cross section for Mueller-Navelet (MN) dijet production at TeV-scale energies [8]. MN dijets refer to the pair of jets in an event that have the largest separation in rapidity among all jets with transverse momentum above a given threshold $p_{\perp\min}$ [37]. The CMS Collaboration also provided comparisons of their measurements with theoretical calculations that incorporate leading order (LO) and next-to-leading order (NLO) QCD matrix elements, enhanced by leading logarithmic (LL) DGLAP evolution of the parton shower. The comparison revealed that theoretical models using DGLAP evolution significantly overestimate the measured dijet production cross sections, by up to a factor of $6$ in the region of large rapidity separation $\Delta y$. Noticeable deviations of DGLAP-based calculations from the measurements begin to appear for rapidity separations $\mbox{$\Delta y$}\gtrsim 4$, indicating that the theoretical models based on the DGLAP evolution, become less accurate in this large rapidity interval region.
The accuracy of LL BFKL-based calculations is generally insufficient, as the LL approximation tends to predict stronger effects than those observed in experimental measurements. This overestimation arises due to limitations inherent in the LL BFKL approach, such as neglecting certain higher-order corrections like running of the strong coupling and kinematical constraints such as energy conservation. Consequently, NLL corrections [20, 21] are necessary to achieve a more accurate and realistic description of the data. The NLL BFKL calculations rely on three key components: the process-independent BFKL Green’s function, process-dependent impact factors, and a procedure to resolve the QCD renormalization scale ambiguity. The detailed calculations for the BFKL Green’s function and impact factors for MN dijets can be found in Refs [38, 39]. The approach that systematically eliminates scale-setting ambiguities was developed by Brodsky, Fadin, Kim, Lipatov and Pivovarov (BFKLP) [40]. This method essentially generalizes the Brodsky, Lepage, and Mackenzie (BLM) optimal scale-setting procedure [41] to the non-Abelian case.
Calculations performed in Ref. [10], based on the NLL BFKL evolution, demonstrate good agreement with the measurement data in the region $\mbox{$\Delta y$}>4$. This provides new evidence supporting the possible manifestation of BFKL evolution effects at TeV energy scales and suggests that the QCD background is more accurately described by the NLL BFKL evolution in the regime of large $\Delta y$.
@@ IV Definition of observable
In the definition of the observable given above, we opted for MN jets instead of inclusive jets due to several reasons.
First, the MN dijet cross section is much easier to compute within the NLL BFKL framework than, for example, the inclusive dijet cross section. Here the inclusive dijet production includes all pairwise jet combinations meeting selection criteria. As shown in Refs. [5] and [8], NLL BFKL calculation better reproduces data at large $\Delta y$ than other QCD approaches.
Second, at large $\Delta y$, the MN dijet cross section approaches the inclusive one. At large $\Delta y$ and $p_{\perp\min}$ thresholds, little energy remains for a third high-$p_{\perp}$ jet with large rapidity. This is even more true when large $M_{jj}$ selections are imposed. In case of large $M_{jj}$ selection nearly all collision energy is spend for the MN dijet production. Additionally, within the DGLAP evolution, leading dijet (leading dijet consists of two jets with highest $p_{\perp}$ in the event) cross sections are more natural to calculate. At large $\Delta y$ and high $M_{jj}$, the MN dijet is almost surely the leading one. Thus, differences between MN, inclusive, and leading dijet cross sections should vanish in this regime.
Third, it is known, the BFKL dynamic produces azimuthal angular decorrelations of jets forming MN dijet due to real QCD emission of additional jets within the rapidity interval between jets [45, 46]. We expect transverse momentum angular correlations of current calculation [employing Eq. (4)] for high-$M_{jj}$ dijets at large $\Delta y$ produced by $t$-channel multiple graviton exchange within ADD gravity to resemble the evolution of DGLAP more than the BFKL one. This is because, the partonic cross section in Eq. (4) omits real graviton emission, so there is no azimuthal decorrelation at the subprocess level. Observing ADD-induced azimuthal decorrelations would require including real graviton emission alongside the sequential graviton exchanges accounted in Eq. (4). This approach, to our knowledge, is not yet successfully implemented. If real graviton emission is included in the calculation in the trans-Planckian eikonal regime, it can produce non-negligible azimuthal decorrelation. From this point of view, the azimuthal decorrelation of jets forming high-mass, large-$\Delta y$ dijets is interesting to measure.
Finally, the measured MN dijet cross section shows more pronounced deviations from the DGLAP predictions [Ref. [8]] than measured azimuthal decorrelations [Ref. [5]]. Specifically, the DGLAP predictions overestimates the measured MN dijet cross section by a factor of 5 at $\mbox{$\Delta y$}\sim 7.5$ for $\sqrt{s}=2.76$ TeV [Ref. [8]], while the deviation of the DGLAP predictions from measured azimuthal decorrelation at $\sqrt{s}=7$ TeV is comparable to experimental systematic uncertainty [Ref. [5]].
This study examines $pp$ collisions at energies of $\mbox{$\sqrt{s}$}=13$, $40$, and $100$ TeV, spanning from current LHC energies up to those expected for FCCpp and CEPC-SppC. Although the ADD signal to QCD background ratio is expected to improve with increasing $M_{jj\mathrm{min}}$, the cross sections diminish significantly at large $M_{jj\mathrm{min}}$. Thus, the high dijet mass cuts are selected based on the projected integrated luminosities of the HL-LHC, FCCpp and CEPC-SppC. Specifically, for $\mbox{$\sqrt{s}$}=13$ TeV we consider $\mbox{$M_{jj\mathrm{min}}$}=6$, and $9$ TeV; for $\mbox{$\sqrt{s}$}=40$ TeV, $\mbox{$M_{jj\mathrm{min}}$}=9$, and $30$ TeV; and for $\mbox{$\sqrt{s}$}=100$ TeV, $\mbox{$M_{jj\mathrm{min}}$}=30$, and $70$ TeV.
@@
$M_{jj}^{2}=2p_{\perp 1}p_{\perp 2}(\cosh{(\mbox{$\Delta y$})}-\cos(\Delta\phi)),$ (5)
where $\Delta\phi$ represents the difference in azimuthal angle between the two jets forming the dijet. From here we see that imposition high $M_{jj\mathrm{min}}$ selection on a dijet effectively select jets with $p_{\perp}$ higher than $p_{\perp\min}$. In the BFKL kinematic regime, $p_{\perp 1}$ can differ from $p_{\perp 2}$ due to transverse momentum diffusion. The effective minimum jet transverse momentum, $p_{\perp\min}^{\mathrm{eff}}$, is determined by $M_{jj\mathrm{min}}$, $\Delta y$ and $p_{\perp\max}=\mbox{$\sqrt{s}$}/(2\cosh(\max[-y_{\max}+\mbox{$\Delta y$},0]))$. For the selections analyzed in this study, the effective minimum jet $p_{\perp}$ in BFKL kinematics is shown in Fig 1(a). It is observed that $\mbox{$p_{\perp\min}^{\mathrm{eff}}$}<\mbox{$p_{\perp\min}$}=20$ GeV only for the $\mbox{$\sqrt{s}$}=40$ TeV with $\mbox{$M_{jj}$}>9$ TeV selection, where $\mbox{$p_{\perp\min}^{\mathrm{eff}}$}\approx 18$ GeV for $\mbox{$\Delta y$}>6$. However, cases with a large imbalance, i.e. when the $p_{\perp}$ of one jet is much smaller than that of the other, lead to an oscillating integrand in the BFKL calculation, causing these configurations to be suppressed relative to those with $p_{\perp 1}\approx p_{\perp 2}$. When $p_{\perp 1}=p_{\perp 2}$, as in the DGLAP kinematics, the effective minimum transverse momentum, $p_{\perp\min}^{\mathrm{eff}}$, is determined by $M_{jj\mathrm{min}}$ and $\Delta y$. The corresponding $p_{\perp\min}^{\mathrm{eff}}$ values are presented in Fig 1(b). It is evident that $p_{\perp\min}^{\mathrm{eff}}$ is significantly higher than $\mbox{$p_{\perp\min}$}=20$ GeV in this case.
@@
V Numerical calculations and systematic uncertainty
@@
$\Delta y>\log\big((R_{\mathrm{S}}M_{jj})^{2}-1\big).$ (6)
In the ADD gravity signal calculation, $M_{jj}$ is integrated from $M_{jj\mathrm{min}}$ to $\sqrt{s}$. Strictly speaking, at fixed $\Delta y$, BH first forms when $\mbox{$M_{jj}$}=\mbox{$\sqrt{s}$}$. However, this is suppressed by the sharp drop in structure functions as $x\rightarrow 1$. For a rough estimate of the lower $\Delta y$ bound, we take $\mbox{$M_{jj}$}=\mbox{$\sqrt{s}$}$. Higher $M_{D}$ at fixed $M_{jj}$ leads to lower $\Delta y$ bound, because gravity becomes weaker, and visa versa. As we are about to see, the most interesting results are for $M_{D}=3$ TeV at $\mbox{$\sqrt{s}$}=13$ TeV; $M_{D}=10$ TeV at $\mbox{$\sqrt{s}$}=40$ TeV; and $M_{D}=20$ TeV at $\mbox{$\sqrt{s}$}=100$ TeV. For these values $\mbox{$\Delta y$}_{\min}$ is found between 3.5 and 5.5 according to Eq. (6). For $\Delta y$ below $\mbox{$\Delta y$}_{\min}$, some signal leakage due to BH formation may occur. Nevertheless, we believe a step toward larger $\Delta y$ values is needed for Eq. (4) to be fully reliable. Thus, we base our conclusions on ADD gravity sensitivity in the trans-Planckian eikonal regime on background-signal comparisons at $\mbox{$\Delta y$}\sim 8.7$. If future measurements reveal an excess over the SM QCD background at lower $\Delta y$, it could still indicate ADD gravity with $M_{D}$ values larger than those considered in our study.
The QCD background within the DGLAP approximation is evaluated by either convolving LO QCD partonic cross sections with NNPDF31_lo_as0130 PDFs evolved through LL DGLAP evolution or using the pythia8 MC generator with the CP5 tune [49]. Originally pythia8 performed at LO+LL DGLAP accuracy as well. However, the CP5 tune updates it by using a NLO strong coupling constant and NNLO PDFs, along with rapidity ordering in the initial-state radiation. This rapidity ordering makes behavior of pythia8 predictions closer to the BFKL kinematics, improving the modeling of gluon emissions and parton evolution for high-energy collisions. The validity of these adjustments require careful theoretical consistency checks. We developed a custom pythia8 plugin to improve computational efficiency in MC simulations by biasing event generation toward high dijet masses at the parton level and correcting this bias with compensating weights to maintain accurate overall distributions. The plugin uses dijet mass at parton level, and its performance was validated against unweighted simulations. The applied exponential weight-bias function enhances computational efficiency of high dijet mass event generation while keeping weights bounded and uncertainties manageable. The form of the bias function is
$\exp\left(\frac{\sqrt{\hat{s}}}{a}\right)-1+\frac{b}{a},$ (7)
@@
The systematic uncertainty of the NLL BFKL calculations are estimated. As previously noted, NLL BFKL calculations depend strongly on renormalization and factorization scales. The BFKLP procedure following the idea of restoring conformal properties of high-energy asymptotic regime fixes optimal value of renormalization and factorization scales tightly. To estimate the scale uncertainty within BFKL calculations it is possible to vary $s_{0}$ parameter, which define the beginning of the high-energy asymptotic. The central value of $s_{0}$ is typically defined by the product $|p_{\perp 1}|\times|p_{\perp 2}|$. It is varied by factors 2 and 0.5 to obtain the uncertainty.
A further source of systematic uncertainty in these calculations stems from incomplete knowledge of the PDFs. To estimate this uncertainty, we use MC replicas from the PDF4LHC15_NLO_MC set [54]. This particular PDF set provides a combination of three PDF sets, namely CT14 [55], MMHT2014 [56], and NNPDF30 [57], and properly accounts for the uncertainties from each of these sets at NLO accuracy.
The total systematic uncertainty of the NLL BFKL calculations combines these contributions in quadrature. When $M_{jj\mathrm{min}}$ is not too close to $\sqrt{s}$, both sources contribute comparably. However, as $M_{jj\mathrm{min}}$ approaches $\sqrt{s}$, the PDF uncertainty dominates due to its growth in the $x\rightarrow 1$ limit.